OpenStudy (anonymous):

Pleaseeee help Find two consecutive odd integers whose sum is 168

OpenStudy (campbell_st):

let x be the 1st odd integer. do you know what the next odd integer would be?

OpenStudy (anonymous):

83+85

OpenStudy (anonymous):

83> 85

OpenStudy (anonymous):

How do you figure it out

OpenStudy (anonymous):

x+(x+2)=168

OpenStudy (anonymous):

x+x+2 = 168 > 2x+2=168 now subtract 2-2 and 168-2 2x=166 2x/2x = 166/2x = 83

OpenStudy (anonymous):

let x be the 1st odd integer let y=x+2

OpenStudy (anonymous):

well that's how i did it, not sure if it's the correct correct way but it sure was correct in the math haha =)

OpenStudy (anonymous):

let x be the 1st int let y be the 2nd int the equation will be: x+y=168 we then let y=x+2 since it is the preceding odd integer. then substitute y in x+y=168,thus x+(x+2)=168. solve.